Separate 152 into 2 parts such that if the bigger number is divided by the smaller number, the result is 5/2.
Found 2 solutions by checkley77, RAY100:
Answer by checkley77(12844)

5/2=X/(152-X)
2X=5*152-5X
2X+5X=760
7X=760
X=760/7
X=108.57142 IS THE LARGER PART.
152-108.57142=43.42858 IS THE SMALLER PART.
PROOF:
108.57142/43.42858~2.5=5/2

Answer by RAY100(1637)

(1) x+y= 152
(2) x/y = 5/2 = 2.5,,,,,,,x=(2.5y)
.
subst (2) into (1)
.
2.5y + y = 152
3.5 y =152
y= 43.43,,,,,,answer
.
subst in (2)
x = 2.5 * 43.43 = 108.57
.
check
(1) 108.57 +43.43 = 152,,,,ok
(2) 108.57 / 43.43 = 2.5,,,,,,ok